OpenStudy (anonymous):

To estimate the height of a tree, Dave stands in the shadow of the tree so that his shadow is and the tree's shadow end at the same point. Dave is 6ft 4inches tall and his shadow is 15ft long. If he is standing 66ft away from the tree, what is the height of the tree?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{x}{81}=\frac{\frac{19}{3}}{15}\] \[\frac{x}{81}=\frac{19}{45}\] \[x=\frac{19\times 81}{15}\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw